Given the components of the two Forces, determine the magnitude and the direction of the two Forces. Tx = 12kN & Ty =-5kN; Ux = -8kN & Uy = 22kN
1
Expert's answer
2021-08-26T08:10:47-0400
"T=\sqrt{12^2+5^2}=13\ kN\\
\theta=360-\arctan{\frac{5}{12}}=337\degree"

"U=\sqrt{8^2+22^2}=23.4\ kN\\
\alpha=180-\arctan{\frac{22}{8}}=110\degree"